# Exploiting Vocabulary Frequency Imbalance in Language Model Pre-training

### v2

@@ 4.1 Frequent-word overlap explains downstream performance transfer
Reducing loss for frequent words is key to achieving a lower global cross-entropy loss. But does this effect carry over to downstream task accuracy? Although a strong link between lower loss and better downstream performance has been studied [20, 24], its rationale has not been investigated. To shed light on this phenomenon, we analyze the overlap of frequent words between the pre-training data and evaluation benchmarks and demonstrate that scaling the vocabulary size reduces cross-entropy loss during pre-training and on downstream tasks. Figure 4(a) shows that the $2,500$ most common words in FineWeb-Edu [40] account for roughly 7676-7878% of all tokens in ARC [3], HellaSwag [63], and SciQ [58] and about 7272% in the PIQA [5] and CC‑Main‑2023402023‑40 dump [21]. Because lower cross‑entropy loss on this CC subset correlates closely with stronger reasoning benchmark scores [21], reducing the frequent words loss and driving down global cross-entropy will improve accuracy across downstream tasks. Consistent with that expectation, figure 4(b) shows that increasing the vocabulary size from $24\mathrm{K}$ to $196\mathrm{K}$ lowers the average per-word loss on the frequent $2,500$ FineWeb-Edu words by about $0.11$ nats, translating into a roughly $0.07$ nats drop in global cross-entropy loss on the CC dataset. Figure 4(c) further confirms that a larger vocabulary also improves average downstream task accuracy of the language model across datasets and model scales. The key insight from this observation is that frequent words highly overlap between the typical training dataset and the downstream benchmark, so the cross‑entropy reduction achieved by enlarging the vocabulary size during pre‑training naturally translates into better downstream performance.
@@
Figure 4: Figure 4(a) demonstrates that the most frequent $2,500$ words in the FineWeb-Edu comprise nearly 727872-78% of the tokens in other downstream benchmark datasets as well as the $C$-Main-2023402023-40 [21]. ARC refers to ARC-Easy, and HS refers to HellaSwag. Figure 4(b) illustrates that a larger vocabulary reduces average per-word loss on frequent FineWeb-Edu words within the CC dataset, and demonstrates how this translates into lower global cross-entropy loss on CC dataset. Figure 4(c) confirms that scaling the vocabulary size boosts downstream task performance.
@@ 4.2 Parameter scaling recovers the same frequent-word gains
Thus far, we have shown that enlarging the vocabulary size decreases word-level average per-word loss on frequent words, which translates into lower cross‑entropy loss and better downstream task performance. One might ask whether similar gains could be achieved without altering the vocabulary size by adjusting other model hyperparameters. Our experiment illustrates that increasing the model’s parameter count can replicate the same benefit. To investigate this, we use the Pythia suite, where each checkpoint is trained on the same dataset with identical hyperparameters, differing only in model size [4]. We also examine the OLMo-2 series [38] to determine whether the same pattern persists in contemporary large-scale language models (see the Appendix D).
Figure 5 compares Pythia models (160160M, 11B, 6.96.9B parameter count) in terms of word-level average per-word loss, proportion of frequent words losses on total loss, and global cross-entropy loss measured on the Paloma validation dataset [35]. Figure 5(a) shows that larger models predict frequent words far more accurately: for the most frequent $2,500$ words in FineWeb-Edu, the average per‑word loss drops from $4.48$ nats (160160M) to $3.76$ nats (11B) and $3.38$ nats (6.96.9B). Furthermore, frequent words loss dominates the total loss as model size grows, which accounts for almost 7070% of the total loss in the Pythia 11B and 6.96.9B model size, compared with roughly 4848% in the Pythia 160160M. Unlike the trend observed when only the vocabulary size is increased (Figure 2(b)), scaling model capacity does not inflate loss on rare tokens; their share of the loss shrinks. Figure 5(b) illustrates that the global cross-entropy loss on the Paloma validation set drops from $4.32$ nats in the Pythia 160160M to $2.51$ nats and $2.26$ nats in the Pythia 11B and 6.96.9B respectively, mirroring the loss reduction seen when expanding the vocabulary (Figure LABEL:fig:embd_b). From this experiment, we can observe that enlarging the model size also lowers loss on frequent words while effectively ignoring rare token losses, so the global cross‑entropy reduction is larger than what scaling up vocabulary size alone can deliver.
@@ 5 Discussion
Table 2: Upper bound of Kolmogorov complexity ($K(X^{N})$) and NCR for the $45.97$ billion‑byte FineWeb‑Edu corpus tokenized with various SuperBPE variants. $t$ and Avg denote the stage-switch vocabulary threshold and an average downstream performance reported by [32], respectively. The SuperBPE variant with the highest average performance exhibits the lowest tokenized text complexity.
superBPE $K(X^{N})$ NCR Avg [32]
@@ 5.1 Can we reduce the tokenized text complexity without intensifying frequency imbalance?
A larger vocabulary reduces the tokenized text complexity by decreasing the number of tokens and sharpening the token-frequency imbalance. However, Section 3.6 shows that a sharper token-frequency imbalance increases rare-token loss. This raises the question: can we lower tokenized text complexity without enlarging the vocabulary size? SuperBPE [32] is one such example.
@@
Table 2 reports the upper bound Kolmogorov complexity, NCR of the $45.97$ billion‑byte FineWeb‑Edu corpus tokenized with various SuperBPE variants, along with average downstream performance reported by [32]. The result shows that the superBPE achieves a lower token count and a less skewed token-frequency distribution than BPE at equal vocabulary. Also, the superBPE variant with the highest average performance exhibits the lowest tokenized text complexity. This finding demonstrates that both vocabulary expansion and the superBPE variants leverage the same benefit: they reduce the complexity of the tokenized text by segmenting common character sequences into single tokens, thereby facilitating the model to learn non-i.i.d. patterns in the data more easily.
@@ 5.2 Deduplication through the lens of vocabulary frequency imbalance
Deduplication removes exact or near-duplicate content from a corpus, reducing repetition and leakage so that a larger fraction of the dataset is novel [28, 40, 31]. Deduplication can increase the information density of a dataset, i.e., the fraction of non-redundant data in the corpus. For non-i.i.d. data, information density can be proxied by the byte-level entropy rate, the theoretical minimum amount of non-redundant information. However, the entropy rate is only tightly approximable with near infinite context compressors, finite window compressors can distort comparisons before and after deduplication. A more practical loose upper bound is the byte-level bits-per-byte (BPB): under an i.i.d. assumption it quantifies information density, with unigram (byte) entropy upper-bounding the true entropy rate. Under the byte-level vocabulary, greater information density implies a higher BPB, driving entropy toward $\log_{2}256=8$ bits/byte, which corresponds to a more uniform distribution and lower frequency imbalance. Exploring how deduplication alters BPB, information density, and vocabulary frequency imbalance is a promising avenue for future work.
@@ 5.3 Why higher token‑frequency imbalance degrades machine translation task performance?
Zouhar et al. [64] reported that higher token-frequency imbalance degrades machine translation task performance, contrary to our findings. What explains this contrasting behavior between machine translation and monolingual settings? We presume that this is related to rare word issues in machine translation tasks [52, 50, 25, 34, 17].
In machine translation, the source and target vocabularies overlap minimally: vocabularies that are common in English often occur rarely in French, and vice versa. Conversely, in a monolingual setting, the pre-training corpus and downstream benchmarks share not only vocabularies but also many of the same frequent words, resulting in extensive text overlap. Consequently, BLEU penalizes every missing target-side n-gram, so mistranslating even a handful of infrequent tokens can sharply lower scores. To address this rare-word challenge, methods like vocabulary trimming deliberately shrink the vocabulary size to reduce the impact of low-frequency tokens on translation performance [15, 43, 8].
@@ 6 Related Work
Empirical and theoretical evidence shows that tokenization is a central determinant of both the quality and the speed with which Transformers learn. Early byte or character-level systems such as CANINE [9] and ByT5 [60] avoid sub‑word vocabularies but pay a steep price: sequences become an order of magnitude longer, gradients are noisier, and convergence is markedly slower than for sub‑word models. Enlarging the BPE dictionary lets even an unigram model approach near-optimal cross-entropy by approximating word-level tokens, whereas the same model without a tokenizer underfits [44]. Furthermore, Controlled scaling studies reveal a log–linear pattern: exponentially expanding the input vocabulary leads to an almost linear drop in loss across model sizes [55]. Huang et al. [20] push this further with “Over‑Tokenized Transformers,” showing that a 400400M parameter model with a 12.812.8M token encoder matches the loss of a 11B model baseline without extra compute.
Beyond the gains from larger vocabularies, a growing body of work shows that permitting merges across word boundaries (i.e., disabling whitespace pretokenization) improves language model performance. Liu et al. [32], Schmidt et al. [49] turn off white-space pretokenization during BPE tokenizer training once the vocabulary reaches $t$, or at frequency-driven transition points. Taken together, these results indicate that without tokenization, language models stall at character granularity, but with a tokenizer, they converge faster and generalize better.
@@
Large language models perform a two stage lossless compression where tokenizer acts as a pre-compressor and transformer models the tokenized data [11, 30]. Lester et al. [30] further explore the role of tokenizer as a pre-compressor by replacing it with small transformer model with arithmetic coding, and models trained on neurally compressed text underperforms subword tokenizers but beat byte-level LMs. Delétang et al. [11] and Valmeekam et al. [57] has been proposed that transformer language models can be harnessed for compression via arithmetic coding and Heurtel-Depeiges et al. [19] extends this line by probing the possibility of transformers as a universal compressor across text, images, and audio. However, most literatures mixuse the concept of generalization and compression where generalization focuses on predicting unseen data from different datasets by exploiting shared statistical patterns while compression focuses on encoding a fixed dataset in as few bits as possible by modeling its distribution and removing redundancy. This work investigates the role and impact of tokenizer on generalization through the lens of data compression and information theory.
@@
BPE tokenizers typically adopt pre-tokenization where regular expressions split text into chunks, sometimes called pre-tokens. Recent work shows that removing pre-tokenization degrades downstream performance across diverse tokenizers [10, 47]. Pre-tokenization rules act as an inductive bias grounded in morphology and grammar, helping the tokenizer capture ubiquitous statistical patterns in natural text, which is crucial for transformer language models to generalize unseen data.
@@ 7 Conclusion
This work set out to explain how and why larger vocabularies boost language model performance. Our experiments reveal a single, robust mechanism: enlarging the vocabulary reduces tokenized text complexity, making non-i.i.d. patterns easier to learn, better approximating natural data’s low intrinsic complexity, and ultimately lowering language modeling difficulty. Once a vocabulary reaches roughly 2424K size, every common word is already a single token; subsequent growth therefore does not refine segmentation but instead steepens the long-tailed frequency distribution, focusing optimization on the same frequent tokens and driving loss down. Norm-constraining that erases the frequency signal removes these gains, confirming causality, while enlarging model parameters with a fixed tokenizer reproduces the benefit, pointing to a shared optimization dynamic between vocabulary and model scaling. Recognising this fact turns Kolmogorov complexity of data into a principled dial for tokenizer–model co-design and sharpens our understanding of the scaling forces that govern language-model pre-training. We therefore conclude:
@@ Appendix A Influence of token-frequency imbalance on unigram and language models
Under a pure unigram model, minimum cross‐entropy loss exactly equals Shannon entropy, and modifying the vocabulary size of the tokenizer immediately changes the Shannon entropy. Typically, enlarging the vocabulary segments frequent multi-token patterns into single tokens, driving their individual relative frequency up and reducing Shannon entropy. But if the vocabulary size grows too large, the new entries tend to be rare tokens, which lengthen the tail and can actually increase the Shannon entropy as token-frequency imbalance rises [64]. However, experimental results show that expanding a BPE vocabulary to around $80\mathrm{K}$ lowers the Shannon entropy of unigram models, demonstrating that a more skewed token-frequency distribution is advantageous at practical vocabulary scales [44]. This pattern extends to n-gram models as well: their conditional Shannon entropy—the lowest possible n-gram cross-entropy—can never exceed the unigram entropy, so lowering the unigram entropy necessarily lowers the conditional entropy.
@@
where $V$ denotes the tokenizer’s vocabulary, $p(x)$ the marginal probability of token $x$, $P(\,\cdot\mid x_{<t})$ the true next-token distribution given the full history $x_{<t}$, and $Q_{\theta}(\,\cdot\mid x_{<t})$ the model’s predicted distribution with parameters $\theta$. When the target label is a one-hot vector, the language model loss can be written as $\mathcal{L}(\theta)=\sum_{x\in V}p(x)\,\bigl[-\log Q_{\theta}(x\mid x_{<t})\bigr]$ so it is a marginal‐frequency‐weighted average of the model’s surprisal $\log Q_{\theta}$. By contrast, the Shannon entropy of the unigram model is a marginal‐frequency‐weighted average of the self-information $\log p(x)$ in the dataset. Even though frequent token logits and embedding norms are higher than rare ones (see Appendix LABEL:apdx:embdnorm and LABEL:apdx:unembd_grad), the loss $\log Q_{\theta}$ depends not only on the underlying relative token-frequencies but also on training dynamics as well. We therefore cannot derive a closed-form expression to predict how much the loss on frequent tokens shrinks versus how much the rare token losses grow under $\mathcal{L}=p_{\mathrm{freq}}\,L_{\mathrm{freq}}+p_{\mathrm{rare}}\,L_{\mathrm{rare}}$. This masking effect makes it substantially harder to measure the influence of token‐frequency imbalance in language models than in the unigram and n-gram models.

### v3

@@ 4.1 Frequent-word overlap explains downstream performance transfer
Reducing loss for frequent words is key to achieving a lower global cross-entropy loss. But does this effect carry over to downstream task accuracy? Although a strong link between lower loss and better downstream performance has been studied [huang2025overtokenizedtransformervocabularygenerally, kim2025perilnrevisitinglayernormalization], its rationale has not been investigated. To shed light on this phenomenon, we analyze the overlap of frequent words between the pre-training data and evaluation benchmarks and demonstrate that scaling the vocabulary size reduces cross-entropy loss during pre-training and on downstream tasks. Figure 4(a) shows that the $2,500$ most common words in FineWeb-Edu [penedo2024finewebdatasetsdecantingweb] account for roughly 7676-7878% of all tokens in ARC [clark2018thinksolvedquestionanswering], HellaSwag [zellers2019hellaswagmachinereallyfinish], and SciQ [sciq] and about 7272% in the PIQA [bisk2019piqareasoningphysicalcommonsense] and CC‑Main‑2023402023‑40 dump [huang2024compressionrepresentsintelligencelinearly]. Because lower cross‑entropy loss on this CC subset correlates closely with stronger reasoning benchmark scores [huang2024compressionrepresentsintelligencelinearly], reducing the frequent words loss and driving down global cross-entropy will improve accuracy across downstream tasks. Consistent with that expectation, figure 4(b) shows that increasing the vocabulary size from $24\mathrm{K}$ to $196\mathrm{K}$ lowers the average per-word loss on the frequent $2,500$ FineWeb-Edu words by about $0.11$ nats, translating into a roughly $0.07$ nats drop in global cross-entropy loss on the CC dataset. Figure 4(c) further confirms that a larger vocabulary also improves average downstream task accuracy of the language model across datasets and model scales. The key insight from this observation is that frequent words highly overlap between the typical training dataset and the downstream benchmark, so the cross‑entropy reduction achieved by enlarging the vocabulary size during pre‑training naturally translates into better downstream performance.
@@
Figure 4: Figure 4(a) demonstrates that the most frequent $2,500$ words in the FineWeb-Edu comprise nearly 727872-78% of the tokens in other downstream benchmark datasets as well as the $C$-Main-2023402023-40 [huang2024compressionrepresentsintelligencelinearly]. ARC refers to ARC-Easy, and HS refers to HellaSwag. Figure 4(b) illustrates that a larger vocabulary reduces average per-word loss on frequent FineWeb-Edu words within the CC dataset, and demonstrates how this translates into lower global cross-entropy loss on CC dataset. Figure 4(c) confirms that scaling the vocabulary size boosts downstream task performance.
@@ 4.2 Parameter scaling recovers the same frequent-word gains
Thus far, we have shown that enlarging the vocabulary size decreases word-level average per-word loss on frequent words, which translates into lower cross‑entropy loss and better downstream task performance. One might ask whether similar gains could be achieved without altering the vocabulary size by adjusting other model hyperparameters. Our experiment illustrates that increasing the model’s parameter count can replicate the same benefit. To investigate this, we use the Pythia suite, where each checkpoint is trained on the same dataset with identical hyperparameters, differing only in model size [biderman2023pythiasuiteanalyzinglarge]. We also examine the OLMo-2 series [olmo2] to determine whether the same pattern persists in contemporary large-scale language models (see the Appendix D).
Figure 5 compares Pythia models (160160M, 11B, 6.96.9B parameter count) in terms of word-level average per-word loss, proportion of frequent words losses on total loss, and global cross-entropy loss measured on the Paloma validation dataset [magnusson2024palomabenchmarkevaluatinglanguage]. Figure 5(a) shows that larger models predict frequent words far more accurately: for the most frequent $2,500$ words in FineWeb-Edu, the average per‑word loss drops from $4.48$ nats (160160M) to $3.76$ nats (11B) and $3.38$ nats (6.96.9B). Furthermore, frequent words loss dominates the total loss as model size grows, which accounts for almost 7070% of the total loss in the Pythia 11B and 6.96.9B model size, compared with roughly 4848% in the Pythia 160160M. Unlike the trend observed when only the vocabulary size is increased (Figure 2(b)), scaling model capacity does not inflate loss on rare tokens; their share of the loss shrinks. Figure 5(b) illustrates that the global cross-entropy loss on the Paloma validation set drops from $4.32$ nats in the Pythia 160160M to $2.51$ nats and $2.26$ nats in the Pythia 11B and 6.96.9B respectively, mirroring the loss reduction seen when expanding the vocabulary (Figure 2 and 3). From this experiment, we can observe that enlarging the model size also lowers loss on frequent words while effectively ignoring rare token losses, so the global cross‑entropy reduction is larger than what scaling up vocabulary size alone can deliver.
@@ 5 Discussion
Table 2: Upper bound of Kolmogorov complexity ($K(X^{N})$) and NCR for the $45.97$ billion‑byte FineWeb‑Edu corpus tokenized with various SuperBPE variants. $t$ and Avg denote the stage-switch vocabulary threshold and an average downstream performance reported by [superbpe], respectively. The SuperBPE variant with the highest average performance exhibits the lowest tokenized text complexity.
superBPE $K(X^{N})$ NCR Avg [superbpe]
@@ 5.1 Can we reduce the tokenized text complexity without intensifying frequency imbalance?
A larger vocabulary reduces the tokenized text complexity by decreasing the number of tokens and sharpening the token-frequency imbalance. However, Section 3.6 shows that a sharper token-frequency imbalance increases rare-token loss. This raises the question: can we lower tokenized text complexity without enlarging the vocabulary size? SuperBPE [superbpe] is one such example.
@@
Table 2 reports the upper bound Kolmogorov complexity, NCR of the $45.97$ billion‑byte FineWeb‑Edu corpus tokenized with various SuperBPE variants, along with average downstream performance reported by [superbpe]. The result shows that the superBPE achieves a lower token count and a less skewed token-frequency distribution than BPE at equal vocabulary. Also, the superBPE variant with the highest average performance exhibits the lowest tokenized text complexity. This finding demonstrates that both vocabulary expansion and the superBPE variants leverage the same benefit: they reduce the complexity of the tokenized text by segmenting common character sequences into single tokens, thereby reducing the number of statistical patterns that models need to learn.
@@ 5.2 Deduplication through the lens of vocabulary frequency imbalance
Deduplication removes exact or near-duplicate content from a corpus, reducing repetition and leakage so that a larger fraction of the dataset is novel [Lee2022, penedo2024finewebdatasetsdecantingweb, Li2024]. Deduplication can increase the information density of a dataset, i.e., the fraction of non-redundant data in the corpus. For non-i.i.d. data, information density can be proxied by the byte-level entropy rate, the theoretical minimum amount of non-redundant information. However, the entropy rate is only tightly approximable with near infinite context compressors, finite window compressors can distort comparisons before and after deduplication. A more practical loose upper bound is the byte-level bits-per-byte (BPB): under an i.i.d. assumption it quantifies information density, with unigram (byte) entropy upper-bounding the true entropy rate. Under the byte-level vocabulary, greater information density implies a higher BPB, driving entropy toward $\log_{2}256=8$ bits/byte, which corresponds to a more uniform distribution and lower frequency imbalance. Exploring how deduplication alters BPB, information density, and vocabulary frequency imbalance is a promising avenue for future work.
@@ 5.3 Why higher token‑frequency imbalance degrades machine translation task performance?
zouhar_tokenization reported that higher token-frequency imbalance degrades machine translation task performance, contrary to our findings. What explains this contrasting behavior between machine translation and monolingual settings? We presume that this is related to rare word issues in machine translation tasks [senrich-rareword, senrich2019, koehn2017-sixchallenge, luong2015, gu-token-level-adaptive].
In machine translation, the source and target vocabularies overlap minimally: vocabularies that are common in English often occur rarely in French, and vice versa. Conversely, in a monolingual setting, the pre-training corpus and downstream benchmarks share not only vocabularies but also many of the same frequent words, resulting in extensive text overlap. Consequently, BLEU penalizes every missing target-side n-gram, so mistranslating even a handful of infrequent tokens can sharply lower scores. To address this rare-word challenge, methods like vocabulary trimming deliberately shrink the vocabulary size to reduce the impact of low-frequency tokens on translation performance [gowda2020, bpe_dropout, bpegetspicky].
@@ 6 Related Work
Empirical and theoretical evidence shows that tokenization is a central determinant of both the quality and the speed with which Transformers learn. Early byte or character-level systems such as CANINE [canine] and ByT5 [byt5] avoid sub‑word vocabularies but pay a steep price: sequences become an order of magnitude longer, gradients are noisier, and convergence is markedly slower than for sub‑word models. Enlarging the BPE dictionary lets even an unigram model approach near-optimal cross-entropy by approximating word-level tokens, whereas the same model without a tokenizer underfits [rajaraman2025theorytokenizationllms]. Furthermore, Controlled scaling studies reveal a log–linear pattern: exponentially expanding the input vocabulary leads to an almost linear drop in loss across model sizes [tao2024scalinglawsvocabularylarger]. huang2025overtokenizedtransformervocabularygenerally push this further with “Over‑Tokenized Transformers,” showing that a 400400M parameter model with a 12.812.8M token encoder matches the loss of a 11B model baseline without extra compute.
Beyond the gains from larger vocabularies, a growing body of work shows that permitting merges across word boundaries (i.e., disabling whitespace pretokenization) improves language model performance. superbpe, schmidt2025boundlessbytepairencoding turn off white-space pretokenization during BPE tokenizer training once the vocabulary reaches $t$, or at frequency-driven transition points. Taken together, these results indicate that without tokenization, language models stall at character granularity, but with a tokenizer, they converge faster and generalize better.
@@
Large language models perform a two stage lossless compression where tokenizer acts as a pre-compressor and transformer models the tokenized data [deletang_modeling, Lester2024]. Lester2024 further explore the role of tokenizer as a pre-compressor by replacing it with small transformer model with arithmetic coding, and models trained on neurally compressed text underperforms subword tokenizers but beat byte-level LMs. deletang_modeling and llmzip has been proposed that transformer language models can be harnessed for compression via arithmetic coding and Depeiges2025 extends this line by probing the possibility of transformers as a universal compressor across text, images, and audio. However, most literatures mixuse the concept of generalization and compression where generalization focuses on predicting unseen data from different datasets by exploiting shared statistical patterns while compression focuses on encoding a fixed dataset in as few bits as possible by modeling its distribution and removing redundancy. This work investigates the role and impact of tokenizer vocabulary size on generalization through the lens of data compression and information theory.
@@
BPE tokenizers typically adopt pre-tokenization where regular expressions split text into chunks, sometimes called pre-tokens. Recent work shows that removing pre-tokenization degrades downstream performance across diverse tokenizers [dagan2024gettingtokenizerpretrainingdomain, schmidt2024]. Pre-tokenization rules act as an inductive bias grounded in morphology and grammar, helping the tokenizer capture ubiquitous statistical patterns in natural text, which is crucial for transformer language models to generalize unseen data.
@@ 7 Conclusion
This work set out to explain how and why larger vocabularies boost language model performance. Our experiments reveal a single, robust mechanism: enlarging the vocabulary reduces tokenized text complexity by segmenting frequent non-i.i.d. patterns into single tokens which in turn decreases the number of statistical patterns the language models learn during pre-training, and ultimately lowering language modeling difficulty. Once a vocabulary reaches roughly 2424K size, every common word is already a single token; subsequent growth therefore minimally refine segmentation but instead steepens the long-tailed frequency distribution, focusing optimization on the same frequent words and driving loss down. Heavy overlap in common words explains downstream performance transfer, while enlarging model parameters with a fixed tokenizer reproduces the benefit, pointing to a shared optimization dynamic between vocabulary and model scaling. Recognising this fact turns Kolmogorov complexity of data into a principled dial for tokenizer–model co-design and sharpens our understanding of the scaling forces that govern language-model pre-training. We therefore conclude:
@@ Appendix A Influence of token-frequency imbalance on unigram and language models
Under a pure unigram model, minimum cross‐entropy loss exactly equals Shannon entropy, and modifying the vocabulary size of the tokenizer immediately changes the Shannon entropy. Typically, enlarging the vocabulary segments frequent multi-token patterns into single tokens, driving their individual relative frequency up and reducing Shannon entropy. But if the vocabulary size grows too large, the new entries tend to be rare tokens, which lengthen the tail and can actually increase the Shannon entropy as token-frequency imbalance rises [zouhar_tokenization]. However, experimental results show that expanding a BPE vocabulary to around $80\mathrm{K}$ lowers the Shannon entropy of unigram models, demonstrating that a more skewed token-frequency distribution is advantageous at practical vocabulary scales [rajaraman2025theorytokenizationllms]. This pattern extends to n-gram models as well: their conditional Shannon entropy—the lowest possible n-gram cross-entropy—can never exceed the unigram entropy, so lowering the unigram entropy necessarily lowers the conditional entropy.
@@
where $V$ denotes the tokenizer’s vocabulary, $p(x)$ the marginal probability of token $x$, $P(\,\cdot\mid x_{<t})$ the true next-token distribution given the full history $x_{<t}$, and $Q_{\theta}(\,\cdot\mid x_{<t})$ the model’s predicted distribution with parameters $\theta$. When the target label is a one-hot vector, the language model loss can be written as $\mathcal{L}(\theta)=\sum_{x\in V}p(x)\,\bigl[-\log Q_{\theta}(x\mid x_{<t})\bigr]$ so it is a marginal‐frequency‐weighted average of the model’s surprisal $\log Q_{\theta}$. By contrast, the Shannon entropy of the unigram model is a marginal‐frequency‐weighted average of the self-information $\log p(x)$ in the dataset. Even though frequent token logits and embedding norms are higher than rare ones, the loss $\log Q_{\theta}$ depends not only on the underlying relative token-frequencies but also on training dynamics as well. We therefore cannot derive a closed-form expression to predict how much the loss on frequent tokens shrinks versus how much the rare token losses grow under $\mathcal{L}=p_{\mathrm{freq}}\,L_{\mathrm{freq}}+p_{\mathrm{rare}}\,L_{\mathrm{rare}}$. This masking effect makes it substantially harder to measure the influence of token‐frequency imbalance in language models than in the unigram and n-gram models.